The length of the tube is 0.9 m
and the temp of the room is 22°C
The first trial the total travel time is 0.004973, the second trial he total travel time is 0.0054054, and the third trial the total travel time is 0.004744, and the fourth trial the time travel is 0.0057106, and the fifth trial the time traveled is 0.0052528, the average is 0.00521716.
what is the speed number and the m/s.
1) Calculate the speed of sound. Remember that your time interval represents the time for sound to travel down the tube and back.
2) The accepted speed of sound at atmospheric pressure and 0C is 331.5 m/s. The speed of sound increases 0.607 m/s for every C. Calculate the speed of sound at the temperature of your room and compare your measured value to the accepted value.
1
Expert's answer
2019-11-18T11:25:35-0500

1) The speed of sound would be "V(1)=L\/t" , where "L=2*0.9=1.8m" - double length of the tube. "t=0.00521716s" - the average time that needed for travel.

"V(1)=1.8\/0.00521716=345.015296m\/s"

2) The speed of sound at atmospheric pressure and 22 oC is "V(2)=331.5+0.607*22=344.854m\/s"

"\Delta V=V(1)-V(2)=345.015-344.854=0.161m/s"